OpenStudy (anonymous):

when sinx=radical3/5, find sin2x and tan2x

OpenStudy (tkhunny):

You will need cos(x). Use the Pythagorean Theorem. \(\sin(2x) = 2\sin(x)\cos(x)\) You will need some other stuff. Let's see if you can find it.

OpenStudy (anonymous):

Okay, using the Pythagorean Theorem, I got cos(x) as 2rad6/25. plugging that into sin(2x)=2sin(x)cos(x) I got 6rad2/25 and I believe that to be incorrect.

OpenStudy (tkhunny):

Is it \(\sin(x) = \dfrac{\sqrt{3}}{5}\) or is it \(\sin(x) = \sqrt{\dfrac{3}{5}}\)